Atom Z, whose atomic mass is 3.49 x 10-22g and the number of its neutrons exceeds the
number of its protons by 41
1
Expert's answer
2020-04-13T02:25:50-0400

Neutrons = Mass number - Atomic Number

Protons = Atomic Number 


Solution:

One mole of Z atoms has a mass equal to the Mass number of Z, and contains 6.02×1023 atoms.

Dividing [Mass number of Z / 6.02×1023] yields a mass of 3.49×10-22 g for one Z atom.

Therefore,

Mass number of Z = 3.49×10-22 * 6.02×1023 = 210.098 amu = 210 amu

According to the assignment: Neutrons = 41 + Protons.

Neutrons = Mass number - Protons.

Then,

41 + Protons = 210 - Protons

2 * Protons = (210-41)

Protons = 169 / 2 = 84.5 = 85

Atom Z = astatine (At).


Answer: Z is astatine (At).